Solve the following linear programming problem using the simplex method:

𝑀𝑎𝑥𝑖𝑚𝑖𝑧𝑒 𝑧 = 14𝑥 + 15𝑦

𝑠𝑢𝑏𝑗𝑒𝑐𝑡 𝑡𝑜

13𝑥 + 15𝑦 ≤ 80

−12𝑥 − 17𝑦 ≥ −120

𝑥 ≥ 0, 𝑦 ≥ 0


1
Expert's answer
2021-09-28T02:21:40-0400

Replace "x" by "x_1" and "y" by "x_2."

Problem is: "\operatorname{Max} Z=14 x_{1}+15 x_{2}"

subject to:

"\begin{array}{rllll} 
& 13 & x_{1} & + & 15 & x_{2} & \leq 80 \\
- & 12 & x_{1} & - & 17 & x_{2} & \leq-120
\end{array}"

Here, "b_{2}=-120<0"

so multiply this constraint by "-1" to make "b_{2}>0 ."

"12 x_{1}+17 \quad x_{2} \geq 120
 
 \ and \ x_{1}, x_{2} \geq 0"

The problem is converted to canonical form by adding slack, surplus and artificial variables as appropriate:

1. As the constraint-1 is of type '≤' we should add slack variable "S_1"


2. As constraint-2 is of type '≥' we should subtract surplus variable "S_2"  and add artificial variable "A_1."


After introducing slack, surplus,artificial variables:


"\operatorname{Max} Z=14 x_{1}+15 x_{2}+0 S_{1}+0 S_{2}-M A_{1}"

subject to:

"13 x_{1}+15 x_{2}+S_{1} \quad=80 \\ 12 x_{1}+17 x_{2} \quad-S_{2}+A_{1}=120 \\ and \ x_{1}, x_{2}, S_{1}, S_{2}, A_{1} \geq 0"




Negative minimum Zj-Cj is -17M-15 and its column index is 2. So, the entering variable is x2.


Minimum ratio is 5.3333 and its row index is 1. So, the leaving basis variable is S1.


∴ The pivot element is 15.


Entering =x2, Departing =S1, Key Element =15






Since all Zj-Cj≥0


Hence, the optimal solution arrives with value of variables as :

x1=0,x2=5.3333


Max Z=80


But this solution is not feasible

because the final solution violates the 2nd constraint  12 x1 + 17 x2 ≥ 120.


and the artificial variable A1 appears in the basis with a positive value of 29.3333